OpenStudy (anonymous):

Find the equation of the line with x-intercept 4 and y-intercept 8. help!!

OpenStudy (anonymous):

y = -2x + 8

OpenStudy (anonymous):

ok can u explain how u got that plzz?

OpenStudy (anonymous):

First we have to find the gradient(m): m = (y2 - y1) / (x2 - x1) m = (0 - 8)/(0 - 4) m = -2 Using the general for finding an eqn of a line: y - y1 = m(x - x1) y - 0 = -2(x - 4) y = -2x + 8